In a class, the average score in a test written by 30 students is 30. If the marks of the top three scorers are excluded the average drops to 28. If the marks scored by the top three students are consecutive integral values, find the score of the topper.
give and in 1hour

Answer by Theo(13342)

if the average for 30 students is equal to 30, then the total score had to be 30 * 30 = 900
if the average for 27 students is equal to 28, then the total score had to be 27 * 28 = 756
the difference is equal to 144.
this is the total score of the 3 students who were taken away.
since their scores are consecutive integral values, and you let x = the score of the lowest scoring person of the three, then the sum of their scores is equal to:
x + x+1 + x+2 = 144
simplify this to get:
3x+3 = 144
subtract 3 from both sides to get:
3x = 141
divide both sides by 3 to get:
x = 47
that's the lowest score.
the three scores are:
47
48
49
the highest score is 49.